If m and k are positive and 10m^2 (k^-1) = 100m, what is m^-1 in terms of k?
Answer by Boreal(15235)

10m^2/k=100m
k=10m^2/100m=(m/10)
so 10k=m
(1/10k)=(1/m)=m^(-1)
-
check
let m=4 so m^(-1)=1/4
So 1/10k=1/4
k=0.4=2/5
10*16(5/2) should equal 100*4=400, which it does.